# Random regular graph states are complex at almost any depth

### v1

@@ 3.3 Proof of Theorem 17
Let us conclude this section by sketching the proof of the random matchings part of LABEL:thm:anticoncentration. The proof proceeds along the same lines as Theorem 17 but is somewhat simpler. We restrict ourselves to outlining the essential difference and leaving the details of adapting the proof as an exercise to the reader.
@@
$\mathbb{E}_{G\sim\mathcal{G}_{m}(n,d),\theta}(m_{2}(G,\theta)$ $=2^{-n}\sum_{L,R\subset[n],L\cap R=\emptyset}\mathbb{E}_{G\sim\mathcal{G}_{m}(n,d)}(-1)^{|A_{G}[L,R]|}$ (104)
@@
and we thus get by Lemmas 20 and 21
$\mathbb{E}_{G\sim\mathcal{G}_{m}(n,d),\theta}(m_{2}(G,\theta)=2^{-n}\!\!\!\!\!\!\!\!\!\!\!\sum_{L,R\subset[n],L\cap R=\emptyset}\left(\sum_{\begin{subarray}{c}i=0\\ i=|L|\mod 2\end{subarray}}^{|L|}K^{n-|L|}_{i}(|R|)\frac{|L|!(n-|L|-i-1)!!}{(|L|-i)!!(n-1)!!}\right)^{d}.$ (106)
@@ 4.1 An exposition of forward and reverse switching
In our proofs in Section 4, we make heavy use of a counting technique called switching, which was introduced by Mckay and Wormald [52]. This is a standard technique in the regular-graph theory literature (and combinatorics more generally), see e.g. [38] for a general discussion. Because we need somewhat precise estimates, and this technique is not very well known in the quantum information literature, we do several key calculations explicitly here (specifically for our problem).
@@
$\mathscr{L}$ =$=\left\{G_{n,d}\mid\varepsilon\subseteq G_{n,d};\varepsilon^{\prime}\subseteq\overline{G}_{n,d}\right\},$ (107)
$\mathscr{M}$ =$=\left\{G_{n,d}\mid\varepsilon\backslash\{uw\}\subseteq G_{n,d},uw\notin G_{n,d};\varepsilon^{\prime}\subseteq\overline{G}_{n,d}\right\}.$ (108)
@@
  • There are at most $(d-1)$ choices such that $w=u_{1}$; that is, the endpoints will not be distinct.
@@
#forbidden $\#\textsf{forbidden cases}=\mathcal{O}\bigg{(}s+(d-1)+(d-1)^{2}+s^{\prime}d\bigg{)}=\mathcal{O}\bigg{(}s+d^{2}+s^{\prime}d\bigg{)}.$ (113)
@@
By a similar argument, the number of choices of $u_{2}w_{2}$ is also given by Equation 114. Hence, the total number of choices of $u_{1}w_{1}$ and $u_{2}w_{2}$ are
@@
#forbidden $\#\textsf{forbidden cases}=\mathcal{O}\bigg{(}s+(d-1)+(d-1)^{2}+s^{\prime}d\bigg{)}=\mathcal{O}\bigg{(}s+d^{2}+s^{\prime}d\bigg{)}.$ (120)
@@ 4.1.1 More examples
$\mathscr{L}$ =$=\left\{G_{n,d}\mid\varepsilon\backslash\{uw\}\subseteq G_{n,d};\varepsilon^{\prime}\backslash\{u^{\prime}w^{\prime}\}\subseteq\overline{G}_{n,d},u^{\prime}w^{\prime}\in{G}_{n,d}\right\},$ (122)
$\mathscr{M}$ =$=\left\{G_{n,d}\mid\varepsilon\backslash\{uw\}\subseteq G_{n,d};\varepsilon^{\prime}\subseteq\overline{G}_{n,d}\right\}$ (123)
using forward and reverse switching. Just as before, given a graph $G\in\mathscr{L}$, we will choose two edges $u_{1}w_{1}$ and $u_{2}w_{2}$ of ${G}\backslash\varepsilon$ such that all six endpoints $u^{\prime},w^{\prime},u_{1},w_{1},u_{2},$ and $w_{2}$ are distinct and, in addition, $w^{\prime}u_{1},w_{1}u_{2},w_{2}u^{\prime}$ are not edges of ${G}$ and not in $\varepsilon^{\prime}$. Then, we will forward switch to get a graph $G^{\prime}$ in $\mathscr{M}$. Similarly, we can reverse switch to go from a graph in $\mathscr{M}$ to a graph in $\mathscr{L}$.
@@ 4.2 Expected number of induced subgraphs
Let $d=n^{c}$, for any choice of constant $10<c<$ 1, and let $\varepsilon,\varepsilon^{\prime}$ be two collections of edges on the vertex set $[n]$, with $\varepsilon\cap\varepsilon^{\prime}=\emptyset$ and $|\varepsilon|=s,|\varepsilon^{\prime}|=s^{\prime}$. Let $uw$ be an edge in $\varepsilon$, and $u^{\prime}w^{\prime}$ be an edge in $\varepsilon^{\prime}$. Then, we have:
@@
$\mathscr{L}_{1}$ =$=\left\{G_{n,d}\mid\varepsilon\backslash\{uw\}\subseteq G_{n,d};\varepsilon^{\prime}\backslash\{u^{\prime}w^{\prime}\}\subseteq\overline{G}_{n,d};uw\in G_{n,d};u^{\prime}w^{\prime}\in\overline{G}_{n,d}\right\},$ (126)
$\mathscr{L}_{2}$ =$=\left\{G_{n,d}\mid\varepsilon\backslash\{uw\}\subseteq G_{n,d};\varepsilon^{\prime}\backslash\{u^{\prime}w^{\prime}\}\subseteq\overline{G}_{n,d};uw\in G_{n,d};u^{\prime}w^{\prime}\notin\overline{G}_{n,d}\right\},$ (127)
$\mathscr{L}_{3}$ =$=\left\{G_{n,d}\mid\varepsilon\backslash\{uw\}\subseteq G_{n,d};\varepsilon^{\prime}\backslash\{u^{\prime}w^{\prime}\}\subseteq\overline{G}_{n,d};uw\notin G_{n,d};u^{\prime}w^{\prime}\in\overline{G}_{n,d}\right\},$ (128)
$\mathscr{L}_{4}$ =$=\left\{G_{n,d}\mid\varepsilon\backslash\{uw\}\subseteq G_{n,d};\varepsilon^{\prime}\backslash\{u^{\prime}w^{\prime}\}\subseteq\overline{G}_{n,d};uw\notin G_{n,d};u^{\prime}w^{\prime}\notin\overline{G}_{n,d}\right\}.$ (129)
@@
$\mathscr{L}_{1}$ =$=\left\{G_{n,d}\mid\varepsilon\subseteq G_{n,d};\varepsilon^{\prime}\subseteq\overline{G}_{n,d}\right\},$ (130)
$\mathscr{L}_{2}$ =$=\left\{G_{n,d}\mid\varepsilon\subseteq G_{n,d};\varepsilon^{\prime}\backslash\{u^{\prime}w^{\prime}\}\subseteq\overline{G}_{n,d},u^{\prime}w^{\prime}\in G_{n,d}\right\},$ (131)
$\mathscr{L}_{3}$ =$=\left\{G_{n,d}\mid\varepsilon\backslash\{uw\}\subseteq G_{n,d},uw\notin G_{n,d};\varepsilon^{\prime}\subseteq\overline{G}_{n,d}\right\},$ (132)
$\mathscr{L}_{4}$ =$=\left\{G_{n,d}\mid\varepsilon\backslash\{uw\}\subseteq G_{n,d},uw\notin G_{n,d};\varepsilon^{\prime}\backslash\{u^{\prime}w^{\prime}\}\subseteq\overline{G}_{n,d},u^{\prime}w^{\prime}\in G_{n,d}\right\}.$ (133)
@@
$\mathscr{L}_{2}\cup\mathscr{L}_{4}$ =$=\left\{G_{n,d}\mid\varepsilon\backslash\{uw\}\subseteq G_{n,d};\varepsilon^{\prime}\backslash\{u^{\prime}w^{\prime}\}\subseteq\overline{G}_{n,d},u^{\prime}w^{\prime}\notin\overline{G}_{n,d}\right\},$ (140)
$\mathscr{L}_{1}\cup\mathscr{L}_{3}$ =$=\left\{G_{n,d}\mid\varepsilon\backslash\{uw\}\subseteq G_{n,d};\varepsilon^{\prime}\subseteq\overline{G}_{n,d}\right\}.$ (141)
Once again, by the forward and reverse switching arguments as in Section 4.1.1, we can show that
@@
Let $d=n^{c}$ for any choice of $10<c<$ 1. Let $\varepsilon$ be a fixed disjoint collection of edges on the vertex set $[n]$ of size $s$ and $uw$ be an edge in $\epsilon$. Then,
$\Pr[\varepsilon\subseteq G_{n,d}]=\left(\frac{n\alpha}{n-d(1-\alpha)}\right)\left(\frac{d}{n}\right)\Pr[\varepsilon\backslash\{uw\}\subseteq G_{n,d}],$ (143)
@@
Given $G\in\mathscr{L}_{0}$, we apply a forward switching operation to get a graph in $\mathscr{L}_{1}$, and given a graph $G\in\mathscr{L}_{1}$, we apply reverse switching to get a graph in $\mathscr{L}_{0}$. By similar arguments as in Section 4.1,
@@
From Equation 148, the lemma follows. ∎
@@
Let $d=n^{c}$ for any choice of $10<c<$ 1. Let $uw$ be an edge. Then,
@@
Follows from Lemma 26 by setting $s=1$. ∎
With Lemma 26 and Lemma 25 we can estimate the probability that we observe an induced subgraph on a fixed set of vertices.
@@
Let $d=n^{c}$ for any choice of $10<c<$ 1. Let $\varepsilon,\varepsilon^{\prime}$ be fixed disjoint collections of edges on the vertex set $[n]$, of size $s,s^{\prime}$ respectively. If we assume that $s\cdot s^{\prime}=o(d)$, then:
@@
Without loss of generality, let $s<s^{\prime}$. By recursively applying Lemma 25, we have that
@@
$=\left(\frac{n\alpha}{n-{d(1-\alpha)}}\right)^{s}\left(\frac{d}{n}\right)^{s}\left(1-\frac{d}{n}\right)^{s}\mathbb{P}\left[\varepsilon^{\prime}\backslash\{u_{1}^{\prime}w_{1}^{\prime},u_{2}^{\prime}w_{2}^{\prime},\ldots,u_{s}^{\prime}w_{s}^{\prime}\}\subseteq\overline{G}_{n,d}\right],$
@@
So, by Corollary 27,
[$\mathbb{P}\left[\{u^{\prime}_{s+1}w^{\prime}_{s+1},\ldots u^{\prime}_{s^{\prime}}w^{\prime}_{s^{\prime}}\}\in\overline{G}_{n,d}\right]=\left(1+\frac{o(d\cdot(s^{\prime}-s))}{n}\right)\left(1-\frac{d}{n}\right)^{s^{\prime}-s}.$ (156)
@@
Let $d=n^{c}$ for any choice of $10<c<$ 1. Let $H$ be a fixed graph with $e$ edges and $v$ vertices with $e\left(\binom{v}{2}-e\right)=o(d)$; $Y_{H}$ denote the number of induced copies of $H$ in $G_{n,d}$ and $\textsf{aut}(H)$ be the number of automorphisms of $H$. Then,
@@
Moreover, there are exactly $\frac{\binom{n}{v}v!}{\text{aut}(H)}$ copies of $H$, and by the linearity of expectation, Equation 158 holds. ∎
@@ 4.3 Expected number of grid graphs
Let $d=n^{c}$ for any choice of $10<c<$ 1. Let $G_{n,d}$ be a random $d$-regular graph on $n$ vertices and let $H$ be a grid graph having $v$ vertices for any $v$ satisfying $v^{2}\cdot d=\mathcal{O}(n)$ and $v=o(d^{1/3})$. Let $Y_{H}$ denote the number of induced copies of $H$ in $G_{n,d}$. Then,
@@
Observe that Equation 163 is greater than $1$ when $d=\omega(\sqrt{n})$.
@@ 4.4 Expected number of sparsified square grid graphs
In the previous section we noted that grid graphs are expected to appear whenever $d=\omega(\sqrt{n})$. We can push this lower bound on the degree further down by considering instead sparsified grid graphs. These are constructed from regular $L\times L$ grid graphs by replacing each edge with $L-1$ vertices and $L$ edges connected in a line. These graphs, on $2L(L-1)^{2}+L^{2}$ vertices, are still universal resource states, as measuring all but one of the qubits on each line in the $Y$ basis, which is equivalent to applying the local complementation operation on that vertex, gives back the $L\times L$ grid graph. Choosing the number of vertices $v=2L(L-1)^{2}+L^{2}$ , the number of edges in a sparsified grid graph is $e=2L^{2}(L-1)$. The number of edges in the complement is
@@
With this information we can reprove Corollary 30 for sparsified grid graphs:
@@
Let $d=n^{c}$ for any choice of $10<c<$ 1. Let $G_{n,d}$ be a random $d$-regular graph on $n$ vertices and let $H$ be a sparsified grid graph having $v$ vertices for any $v$ satisfying $v^{2}\cdot d=\mathcal{O}(n)$ and $v=o(d^{1/3})$. Let $Y_{H}$ denote the number of induced copies of $H$ in $G_{n,d}$. Then,

### v2

@@ 3.3 Proof of Theorem˜17
Let us conclude this section by sketching the proof of the random matchings part of LABEL:thm:anticoncentration. The proof proceeds along the same lines as Theorem˜17 but is somewhat simpler. We restrict ourselves to outlining the essential difference and leaving the details of adapting the proof as an exercise to the reader.
@@
$\hskip-4.30554pt\mathbb{E}_{G\sim\mathcal{G}_{m}(n,d),\theta}(m_{2}(G,\theta)$ $=2^{-n}\sum_{L,R\subset[n],L\cap R=\emptyset}\mathbb{E}_{G\sim\mathcal{G}_{m}(n,d)}(-1)^{|A_{G}[L,R]|}$ (104)
@@
and we thus get by Lemmas˜20 and 21
$\mathbb{E}_{G\sim\mathcal{G}_{m}(n,d),\theta}(m_{2}(G,\theta)=2^{-n}\!\!\!\!\!\!\!\!\!\!\!\sum_{L,R\subset[n],L\cap R=\emptyset}\left(\sum_{\begin{subarray}{c}i=0\\ i=|L|\mod 2\end{subarray}}^{|L|}\hskip-8.00003ptK^{n-|L|}_{i}(|R|)\frac{|L|!(n-|L|-i-1)!!}{(|L|-i)!!(n-1)!!}\right)^{d}.\hskip-8.00003pt$ (106)
@@ 4.1 An exposition of forward and reverse switching
In our proofs in Section˜4, we make heavy use of a counting technique called switching, which was introduced by Mckay and Wormald [71]. This is a standard technique in the regular-graph theory literature (and combinatorics more generally), see e.g. [53] for a general discussion. Because we need somewhat precise estimates, and this technique is not very well known in the quantum information literature, we do several key calculations explicitly here (specifically for our problem).
@@
$\mathscr{L}$ $=\left\{G_{n,d}\mid\varepsilon\subseteq G_{n,d};\varepsilon^{\prime}\subseteq\overline{G}_{n,d}\right\},$ (107)
$\mathscr{M}$ $=\left\{G_{n,d}\mid\varepsilon\backslash\{uw\}\subseteq G_{n,d},uw\notin G_{n,d};\varepsilon^{\prime}\subseteq\overline{G}_{n,d}\right\}.$ (108)
@@
  • There are at most (d1)(d-1) choices such that $w=u_{1}$; that is, the endpoints will not be distinct.
@@
$\textsf{forbidden cases}=\mathcal{O}\bigg(s+(d-1)+(d-1)^{2}+s^{\prime}d\bigg)=\mathcal{O}\bigg(s+d^{2}+s^{\prime}d\bigg).$ (113)
@@
By a similar argument, the number of choices of $u_{2}w_{2}$ is also given by Equation˜114. Hence, the total number of choices of $u_{1}w_{1}$ and $u_{2}w_{2}$ are
@@
$\textsf{forbidden cases}=\mathcal{O}\bigg(s+(d-1)+(d-1)^{2}+s^{\prime}d\bigg)=\mathcal{O}\bigg(s+d^{2}+s^{\prime}d\bigg).$ (120)
@@ 4.1.1 More examples
$\mathscr{L}$ $=\left\{G_{n,d}\mid\varepsilon\backslash\{uw\}\subseteq G_{n,d};\varepsilon^{\prime}\backslash\{u^{\prime}w^{\prime}\}\subseteq\overline{G}_{n,d},u^{\prime}w^{\prime}\in{G}_{n,d}\right\},$ (122)
$\mathscr{M}$ $=\left\{G_{n,d}\mid\varepsilon\backslash\{uw\}\subseteq G_{n,d};\varepsilon^{\prime}\subseteq\overline{G}_{n,d}\right\}$ (123)
using forward and reverse switching. Just as before, given a graph $G\in\mathscr{L}$, we will choose two edges $u_{1}w_{1}$ and $u_{2}w_{2}$ of ${G}\backslash\varepsilon$ such that all six endpoints $u^{\prime},w^{\prime},u_{1},w_{1},u_{2},$ and $w_{2}$ are distinct and, in addition, $w^{\prime}u_{1},w_{1}u_{2},w_{2}u^{\prime}$ are not edges of G{G} and not in $\varepsilon^{\prime}$. Then, we will forward switch to get a graph $G^{\prime}$ in $\mathscr{M}$. Similarly, we can reverse switch to go from a graph in $\mathscr{M}$ to a graph in $\mathscr{L}$.
@@ 4.2 Expected number of induced subgraphs
Let $d=n^{c}$, for any choice of constant $0<c<1$, and let $\varepsilon,\varepsilon^{\prime}$ be two collections of edges on the vertex set $[n]$, with $\varepsilon\cap\varepsilon^{\prime}=\emptyset$ and $|\varepsilon|=s,|\varepsilon^{\prime}|=s^{\prime}$. Let $uw$ be an edge in $\varepsilon$, and $u^{\prime}w^{\prime}$ be an edge in $\varepsilon^{\prime}$. Then, we have:
@@
$\mathscr{L}_{1}$ $=\left\{G_{n,d}\mid\varepsilon\backslash\{uw\}\subseteq G_{n,d};\varepsilon^{\prime}\backslash\{u^{\prime}w^{\prime}\}\subseteq\overline{G}_{n,d};uw\in G_{n,d};u^{\prime}w^{\prime}\in\overline{G}_{n,d}\right\},$ (126)
$\mathscr{L}_{2}$ $=\left\{G_{n,d}\mid\varepsilon\backslash\{uw\}\subseteq G_{n,d};\varepsilon^{\prime}\backslash\{u^{\prime}w^{\prime}\}\subseteq\overline{G}_{n,d};uw\in G_{n,d};u^{\prime}w^{\prime}\notin\overline{G}_{n,d}\right\},$ (127)
$\mathscr{L}_{3}$ $=\left\{G_{n,d}\mid\varepsilon\backslash\{uw\}\subseteq G_{n,d};\varepsilon^{\prime}\backslash\{u^{\prime}w^{\prime}\}\subseteq\overline{G}_{n,d};uw\notin G_{n,d};u^{\prime}w^{\prime}\in\overline{G}_{n,d}\right\},$ (128)
$\mathscr{L}_{4}$ $=\left\{G_{n,d}\mid\varepsilon\backslash\{uw\}\subseteq G_{n,d};\varepsilon^{\prime}\backslash\{u^{\prime}w^{\prime}\}\subseteq\overline{G}_{n,d};uw\notin G_{n,d};u^{\prime}w^{\prime}\notin\overline{G}_{n,d}\right\}.$ (129)
@@
$\mathscr{L}_{1}$ $=\left\{G_{n,d}\mid\varepsilon\subseteq G_{n,d};\varepsilon^{\prime}\subseteq\overline{G}_{n,d}\right\},$ (130)
[…]
$\mathscr{L}_{3}$ $=\left\{G_{n,d}\mid\varepsilon\backslash\{uw\}\subseteq G_{n,d},uw\notin G_{n,d};\varepsilon^{\prime}\subseteq\overline{G}_{n,d}\right\},$ (132)
$\mathscr{L}_{4}$ $=\left\{G_{n,d}\mid\varepsilon\backslash\{uw\}\subseteq G_{n,d},uw\notin G_{n,d};\varepsilon^{\prime}\backslash\{u^{\prime}w^{\prime}\}\subseteq\overline{G}_{n,d},u^{\prime}w^{\prime}\in G_{n,d}\right\}.$ (133)
@@
$\mathscr{L}_{2}\cup\mathscr{L}_{4}$ $=\left\{G_{n,d}\mid\varepsilon\backslash\{uw\}\subseteq G_{n,d};\varepsilon^{\prime}\backslash\{u^{\prime}w^{\prime}\}\subseteq\overline{G}_{n,d},u^{\prime}w^{\prime}\notin\overline{G}_{n,d}\right\},$ (140)
$\mathscr{L}_{1}\cup\mathscr{L}_{3}$ $=\left\{G_{n,d}\mid\varepsilon\backslash\{uw\}\subseteq G_{n,d};\varepsilon^{\prime}\subseteq\overline{G}_{n,d}\right\}.$ (141)
Once again, by the forward and reverse switching arguments as in Section˜4.1.1, we can show that
@@
Let $d=n^{c}$ for any choice of $0<c<1$. Let $\varepsilon$ be a fixed disjoint collection of edges on the vertex set $[n]$ of size $s$ and $uw$ be an edge in $\epsilon$. Then,
$Pr[\varepsilon\subseteq G_{n,d}]=\left(\frac{n\alpha}{n-d(1-\alpha)}\right)\left(\frac{d}{n}\right)\Pr[\varepsilon\backslash\{uw\}\subseteq G_{n,d}],$ (143)
@@
Given $G\in\mathscr{L}_{0}$, we apply a forward switching operation to get a graph in $\mathscr{L}_{1}$, and given a graph $G\in\mathscr{L}_{1}$, we apply reverse switching to get a graph in $\mathscr{L}_{0}$. By similar arguments as in Section˜4.1,
@@
From Equation˜148, the lemma follows. ∎
@@
Let $d=n^{c}$ for any choice of $0<c<1$. Let $uw$ be an edge. Then,
@@
Follows from Lemma˜26 by setting $s=1$. ∎
With Lemma˜26 and Lemma˜25 we can estimate the probability that we observe an induced subgraph on a fixed set of vertices.
@@
Let $d=n^{c}$ for any choice of $0<c<1$. Let $\varepsilon,\varepsilon^{\prime}$ be fixed disjoint collections of edges on the vertex set $[n]$, of size $s,s^{\prime}$ respectively. If we assume that $s\cdot s^{\prime}=o(d)$, then:
@@
Without loss of generality, let $s<s^{\prime}$. By recursively applying Lemma˜25, we have that
@@
[…]
@@
So, by Corollary˜27,
$\mathbb{P}\left[\{u^{\prime}_{s+1}w^{\prime}_{s+1},\ldots u^{\prime}_{s^{\prime}}w^{\prime}_{s^{\prime}}\}\in\overline{G}_{n,d}\right]=\left(1+\frac{o(d\cdot(s^{\prime}-s))}{n}\right)\left(1-\frac{d}{n}\right)^{s^{\prime}-s}.$ (156)
@@
Let $d=n^{c}$ for any choice of $0<c<1$. Let $H$ be a fixed graph with $e$ edges and $v$ vertices with $e\left(\binom{v}{2}-e\right)=o(d)$; $Y_{H}$ denote the number of induced copies of $H$ in $G_{n,d}$ and $\textsf{aut}(H)$ be the number of automorphisms of $H$. Then,
@@
Moreover, there are exactly $\frac{\binom{n}{v}v!}{\text{aut}(H)}$ copies of $H$, and by the linearity of expectation, Equation˜158 holds. ∎
@@ 4.3 Expected number of grid graphs
Let $d=n^{c}$ for any choice of $0<c<1$. Let $G_{n,d}$ be a random $d$-regular graph on $n$ vertices and let $H$ be a grid graph having $v$ vertices for any $v$ satisfying $v^{2}\cdot d=\mathcal{O}(n)$ and $v=o(d^{1/3})$. Let $Y_{H}$ denote the number of induced copies of $H$ in $G_{n,d}$. Then,
@@
Observe that Equation˜163 is greater than $1$ when $d=\omega(\sqrt{n})$.
@@ 4.4 Expected number of sparsified square grid graphs
In the previous section we noted that grid graphs are expected to appear whenever $d=\omega(\sqrt{n})$. We can push this lower bound on the degree further down by considering instead sparsified grid graphs. These are constructed from regular $L\times L$ grid graphs by replacing each edge with L1L-1 vertices and $L$ edges connected in a line. These graphs, on $2L(L-1)^{2}+L^{2}$ vertices, are still universal resource states, as measuring all but one of the qubits on each line in the $Y$ basis, which is equivalent to applying the local complementation operation on that vertex, gives back the $L\times L$ grid graph. Choosing the number of vertices $v=2L(L-1)^{2}+L^{2}$ , the number of edges in a sparsified grid graph is $e=2L^{2}(L-1)$. The number of edges in the complement is
@@
With this information we can reprove Corollary˜30 for sparsified grid graphs:
@@
Let $d=n^{c}$ for any choice of $0<c<1$. Let $G_{n,d}$ be a random $d$-regular graph on $n$ vertices and let $H$ be a sparsified grid graph having $v$ vertices for any $v$ satisfying $v^{2}\cdot d=\mathcal{O}(n)$ and $v=o(d^{1/3})$. Let $Y_{H}$ denote the number of induced copies of $H$ in $G_{n,d}$. Then,